# Khoroshkin–Tolstoy approach for quantum superalgebras

### v2

@@ Appendix B Inverse of $B_{q}$
In this appendix we find the inverse of the matrix $B_{q}=([B_{ij}]_{q})_{i,j=1}^{M+N}$. The matrix $B_{q}$ is tridiagonal and to find its inverse we use the results of the paper [55]. Let $U=(u_{ij})_{i,j=1}^{L}$ be an arbitrary tridiagonal matrix. Denote $\alpha_{i}=U_{i,\,i-1}$, $i=2,\ldots,L$, $\beta_{i}=U_{ii}$, $i=1,\ldots,L$, and $\gamma_{i}=U_{i,\,i+1}$, $i=1,\ldots,L-1$. Introduce the following notation for the principal minors of $U$,
$\Theta_{k}=\det(U_{ij})_{i,j=1}^{k},\qquad\Phi_{k}=\det(U_{ij})_{i,j=k}^{N}.$
@@
$\Theta_{-1}=0,\qquad\Theta_{0}=1,\qquad\Phi_{L+1}=1,\qquad\Phi_{L+2}=0,$
@@
$\Phi_{i}=\beta_{i}\Phi_{i+1}-\gamma_{i}\alpha_{i+1}\Phi_{i+2},$ (B.2)
@@
$U^{-1}_{ij}=\left\{\begin{array}[]{ll}(-1)^{i+j}\gamma_{i}\gamma_{i+1}\ldots\gamma_{j-1}\Theta_{i-1}\Phi_{j+1}/\Theta_{N},&i<j,\\[5.0pt] \Theta_{i-1}\Phi_{i+1}/\Theta_{N},&i=j,\\[5.0pt] (-1)^{i+j}\alpha_{j+1}\alpha_{j+2}\ldots\alpha_{i}\Theta_{j-1}\Phi_{i+1}/\Theta_{N},&i>j.\end{array}\right.$ (B.3)
The matrix $B_{q}$ is symmetric, therefore it is enough to find matrix elements $B_{q}^{-1}$ only for $i\leq j$. We have
@@
$\Phi_{i}=\left\{\begin{array}[]{ll}(-1)^{N}[M-N-i+1]_{q},&i\leq M,\\[5.0pt] (-1)^{M+N-i}[M+N-i+1]_{q},&i>M.\end{array}\right.$
@@
$(B_{q}^{-1})_{ij}=\left\{\begin{array}[]{ll}\hskip 8.99994pt[i]_{q}[M-N-j]_{q}/[M-N]_{q},&j<M,\\[5.0pt] -[i]_{q}[N]_{q}/[M-N]_{q},&j=M,\\[5.0pt] -[i]_{q}[M+N-j]_{q}/[M-N]_{q},&i<M,\quad j>M,\\[5.0pt] \hskip 8.99994pt[M]_{q}[M+N-j]_{q}/[M-N]_{q},&i=M,\quad j>M,\\[5.0pt] -[2M-i]_{q}[M+N-j]_{q}/[M-N]_{q}&i>M\end{array}\right.$ (B.4)

### v3

@@ Appendix B Inverse of $B_{q}$
In this appendix we find an explicit expression for the matrix entries of the matrix $C_{q}=B_{q}^{-1}$, where the matrix $B_{q}$ is defined by equation (8.25). The matrix $B_{q}$ is tridiagonal and to find its inverse we use the results of the paper [56]. Let $M=(m_{ij})_{i,j=1}^{L}$ be an arbitrary tridiagonal matrix of size $L$. Introduce the notation $\alpha_{i}=m_{i,\,i-1}$, $i=2,\ldots,L$, $\beta_{i}=m_{ii}$, $i=1,\ldots,L$, and $\gamma_{i}=m_{i,\,i+1}$, $i=1,\ldots,L-1$. Denote also the principal minors of $M$ as
$\theta_{k}=\det(m_{ij})_{i,j=1}^{k},\qquad\varphi_{k}=\det(m_{ij})_{i,j=k}^{L}.$
@@
$\theta_{-1}=0,\qquad\theta_{0}=1,\qquad\varphi_{L+1}=1,\qquad\varphi_{L+2}=0,$
@@
$\varphi_{i}=\beta_{i}\varphi_{i+1}-\gamma_{i}\alpha_{i+1}\varphi_{i+2},$ (B.2)
@@
$m^{-1}_{ij}=\left\{\begin{array}[]{ll}(-1)^{i+j}\gamma_{i}\gamma_{i+1}\ldots\gamma_{j-1}\theta_{i-1}\varphi_{j+1}/\theta_{N},&i<j,\\[5.0pt] \theta_{i-1}\varphi_{i+1}/\theta_{N},&i=j,\\[5.0pt] (-1)^{i+j}\alpha_{j+1}\alpha_{j+2}\ldots\alpha_{i}\theta_{j-1}\varphi_{i+1}/\theta_{N},&i>j.\end{array}\right.$ (B.3)
The matrix $B_{q}$ is symmetric, therefore it is enough to find the matrix elements $B_{q}^{-1}$ only for $i\leq j$. We have
@@
$\varphi_{i}=\left\{\begin{array}[]{ll}(-1)^{N}[M-N-i+1]_{q},&i\leq M,\\[5.0pt] (-1)^{M+N-i}[M+N-i+1]_{q},&i>M.\end{array}\right.$
@@
$c_{qij}=\left\{\begin{array}[]{ll}\hskip 8.99994pt[i]_{q}[M-N-j]_{q}/[M-N]_{q},&j<M,\\[5.0pt] -[i]_{q}[N]_{q}/[M-N]_{q},&j=M,\\[5.0pt] -[i]_{q}[M+N-j]_{q}/[M-N]_{q},&i<M,\quad j>M,\\[5.0pt] -[M]_{q}[M+N-j]_{q}/[M-N]_{q},&i=M,\quad j>M,\\[5.0pt] -[2M-i]_{q}[M+N-j]_{q}/[M-N]_{q}&i>M\end{array}\right.$ (B.4)